Find: 3 whole ½ ÷ (8/3) Find: 3 whole ½ ÷ (8/3) Find: 3 whole ½ ÷ (8/3)

Find:
3 whole ½ ÷ (8/3)
Find:
3 whole ½ ÷ (8/3)
Find:
3 whole ½ ÷ (8/3)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

First covert the mixed fraction into improper fraction.
We have,
= 3 whole 1⁄2
= 7/2
Then,
= (7/2) × reciprocal of (8/3)
= (7/2) × (3/8)
= (7 × 3) / (2 × 8)
= 21/16

Thanks